### v1

# Knowledge Distillation with Structured Chain-of-Thought for Text-to-SQL

Abstract
Deploying accurate Text-to-SQL systems at the enterprise level faces a difficult trilemma involving cost, security and performance. Current solutions force enterprises to choose between expensive, proprietary Large Language Models (LLMs) and low-performing Small Language Models (SLMs). Efforts to improve SLMs often rely on distilling reasoning from large LLMs using unstructured Chain-of-Thought (CoT) traces, a process that remains inherently ambiguous. Instead, we hypothesize that a formal, structured reasoning representation provides a clearer, more reliable teaching signal, as the Text-to-SQL task requires explicit and precise logical steps. To evaluate this hypothesis, we propose Struct-SQL, a novel Knowledge Distillation (KD) framework that trains an SLM to emulate a powerful large LLM. Consequently, we adopt a query execution plan as a formal blueprint to derive this structured reasoning. Our SLM, distilled with structured CoT, achieves an absolute improvement of 8.1% over an unstructured CoT distillation baseline. A detailed error analysis reveals that a key factor in this gain is a marked reduction in syntactic errors. This demonstrates that teaching a model to reason using a structured logical blueprint is beneficial for reliable SQL generation in SLMs.
1 Introduction
Text-to-SQL (NL2SQL or text2sql) has the potential to democratize data access [fei2014constructing]. The field has seen substantial performance advancements driven by the advent of Large Language Models (LLMs) [shi2025asurvey]. These models enhance the capabilities of natural language interfaces for databases by automatically translating natural-language user questions into SQL queries [hong2024next]. Nevertheless, widespread adoption in enterprises remains challenging due to a difficult trade-off among three interdependent factors: cost, security, and performance. This challenge can be understood as an Adoption Trilemma.
Cost: High-performing models typically require significant computational resources, leading to high operational costs, whether using proprietary APIs or privately hosted LLMs [shi2025asurvey].
Security: Relying on external APIs raises prominent security concerns, as transmitting potentially sensitive database schemas and sample records to third-party providers is often unacceptable in enterprise settings [hoang-etal-2025-distill].
Performance: Selecting open-source models for local, private deployment to address cost and security issues often leads to the use of Small Language Models (SLMs), which typically lack adequate zero-shot accuracy for complex real-world queries [li2024can].
To identify the limitations of current approaches, we examine the impact of recent advances in reasoning-driven prompting on Text-to-SQL performance. A notable portion of the recent performance gain in Text-to-SQL using LLMs can be attributed to in-context learning (ICL) [dong-etal-2024-survey]. ICL-based methods, particularly those that use decomposition and multi-step reasoning, have demonstrated substantial gains in performance on the Text-to-SQL task [hong2024next]. A prominent ICL technique for this is Chain-of-Thought (CoT), which encourages models to think step-by-step [wei2022chain, liu2025uncovering]. Such decomposition strategies are beneficial for generating complex SQL queries. For instance, DAIL-SQL [gao2024dail] enhances CoT through context-aware examples, while DIN-SQL [pourreza2023dinsql] and Divide-and-Conquer CoT (DC-CoT) [pourreza2024chase] improve accuracy by breaking questions into intermediate sub-queries. Building on this logical foundation, the Query Plan CoT (QP-CoT) guides the model through steps that mirror a database execution plan, simulating the logical process by which databases execute queries. By following this structured path, QP-CoT ensures that the SQL generation path aligns with the inherent logic of query execution [pourreza2024chase]. Although these reasoning techniques achieve considerable success, their effectiveness is observed almost exclusively with large LLMs. This reliance on large LLMs constitutes a significant limitation: these methods depend on the very models that exacerbate the cost and security challenges, thereby limiting their applicability in resource-constrained enterprise settings and contributing to the ‘Adoption Trilemma’.
The performance limitations of SLMs, suitable for private deployment, are particularly pronounced in the Text-to-SQL domain. Standard benchmarks highlight a significant gap between large models and smaller open-source alternatives. For example, on the BIRD mini-dev benchmark, LLMs such as GPT-4o and Claude (150B parameters) achieve execution accuracies between 30% and 45%, whereas widely used SLMs such as Mistral, Mixtral and Qwen2.5 Coder (7B parameters) achieve only 4% to 12%111Performance on BIRD mini-dev, see https://github.com/bird-bench/mini_dev. This paper reveals that this severe performance degradation in SLMs persists even when advanced reasoning prompts are employed, despite their effectiveness in larger LLMs. This finding is consistent with recent literature, suggesting that large LLMs can adhere to the schema as long as the schema fits within their context window, a capability not observed in SLMs [maamari2024death]. In our experiments, the SLM (Qwen3-4B-Instruct-2507) prompted with QP-CoT fails primarily due to inadequate schema adherence, with a pronounced tendency toward schema hallucination, often generating non-existent tables or columns (see Figure 3(b)). Thus, while large LLMs benefit from structured reasoning strategies, SLMs do not internalize these logical decompositions. This breakdown in structured reasoning motivates an investigation into whether the structure of reasoning itself can be effectively transferred from large LLMs to SLMs.
To bridge this performance gap, Knowledge Distillation (KD) [hinton2014distilling], which transfers reasoning ability from a capable teacher model, is a key strategy [shi2025asurvey]. KD is a model compression technique in which a smaller "student" model is trained to mimic the behavior of a larger, pretrained "teacher" model. Beyond compression, KD enables the transfer of complex reasoning [li2024explanations] and logical skills [hsieh2023cotdistilling, yao2023specializing]. Through KD, our aim is to build SLMs for Text-to-SQL that deliver accuracy comparable to larger LLMs while meeting enterprise cost and security requirements through private deployment. This work argues that KD is a promising approach in addressing this trilemma, but its effectiveness depends on the type of reasoning being distilled.
Given that the core bottleneck for SLMs is replicating the teacher’s complex reasoning, the central question for KD is identifying the optimal form of the reasoning signal to distill. Building on the success of CoT prompting [wei2022chain], one approach is to distill the teacher’s natural language reasoning traces [li2024explanations]. This approach is exemplified by ReasonSQL, where the student model is trained on the teacher’s intermediate CoT steps in addition to the final SQL query. This unstructured approach provides a better knowledge signal than finetuning on SQL queries alone and has been shown to improve model accuracy [rossiello2025rationalization]. However, we argue that the structure of the reasoning signal itself is critical for effective distillation. We hypothesize that distilling knowledge using a more formal, structured representation of the reasoning process that directly reflects the logical steps of query execution could provide a clearer, less ambiguous, and more beneficial supervisory signal for distilling SLMs than unstructured CoT explanations.
To evaluate this hypothesis, we introduce Struct-SQL, a framework for distilling structured reasoning. Figure 1 shows the KD workflow, contrasting the proposed structured reasoning approach against the standard unstructured CoT method. Within this framework, a state-of-the-art (SOTA) Teacher Model is used to generate a QP-CoT trace, which formally decomposes the query into a logical execution plan [pourreza2024chase]. This structured plan, together with the generated SQL query, constitutes the supervisory signal. A student model is then trained to replicate the entire structured output sequence (query plan, SQL). The formal query plan serves as a clear, hierarchical blueprint that guides the student model in learning the precise, logical steps of query construction, from schema linking and join-path selection to aggregation and filtering. During inference, this structured reasoning is retained: the student model is given the QP-CoT prompt to autonomously generate the query plan before synthesizing the final SQL query.
We perform an extensive comparative analysis on the mini-dev BIRD benchmark [li2024can], evaluating our Struct-SQL framework against key baselines. Our contributions are as follows.
We are the first to systematically evaluate the impact of KD using a structured reasoning signal for Text-to-SQL.
We provide a comprehensive error analysis, showing that the improvement in Struct-SQL results from a marked reduction in syntactic errors (e.g., schema hallucination), demonstrating that the structured signal provides a clearer curriculum.
We validate the generalization of our framework through an ablation study on two different SLMs.
We release the KD dataset of 1,300 structured reasoning traces, the Struct-SQL model, prompts and the code 222https://github.com/craterlabs/Struct-SQL-Distillation/.
Following this introduction, Section 2 details the methodology. Section 3 presents the results, which are contextualized by the related work in Section 4, and the paper concludes in Section 5.
2 Methodology
2.1 Problem Formulation
The standard Text-to-SQL task involves mapping a natural language question and a database schema to an executable SQL query . Formally, given an input pair , the goal is to learn a model such that , where represents the predicted query intended to give the same result as the ground truth . Our work focuses on transferring the capabilities of a large, high-performance Teacher Model, , to a smaller, more efficient Student Model, , through KD. We formulate the distillation task as learning the Teacher’s intermediate reasoning steps, , in addition to its final output, . Let represent the complete output sequence. Training is conducted over a distillation dataset . We employ a standard sequence completion loss, minimizing the negative log-likelihood of the teacher-generated text by updating the parameters of the student model:
We instantiate either as an unstructured CoT or a structured QP-CoT to test the hypothesis that a structured reasoning trace provides a more effective supervisory signal.
2.2 Structured CoT via Query Execution Plan
Our structured CoT strategy is inspired by a database engine’s query execution plan. A query execution plan defines the precise sequence of steps that a database follows to access and manipulate data, often generated using an EXPLAIN command. We adopt this ICL prompt strategy from a recent work on a Query Execution Plan-based prompting strategy, QP-CoT [pourreza2024chase]. Instead of a free-form explanation, the Teacher Model is prompted to generate a query plan along with the final SQL query. As shown in Figure 2, this plan decomposes the query into a sequential execution flow that explicitly performs selections, filters, joins, and groupings through step-by-step table scanning and data manipulation, providing a systematic signal for the student model. Although the query plan has been explored as a prompting technique [pourreza2024chase], the key contribution here is to demonstrate the effectiveness of this structured reasoning as the primary teaching signal within a KD framework.
2.3 Experimental Setup
The experiments use the SQLite-based BIRD benchmark for the data [li2024can]. The BIRD training dataset is used for model training, while the BIRD mini-dev dataset serves for evaluation. The Teacher Model (), GPT-4o (OpenAI), is selected due to its leading performance on the BIRD mini-dev dataset and functions as an oracle to generate high-quality query plans. The Student Model () is Qwen3-4B-Instruct-2507 (Alibaba Cloud). It was chosen for its strong performance-to-size ratio, which enables low-latency, private deployment and thus addresses the Adoption Trilemma [yang2025qwen3]. All models utilize a single-pass inference mechanism and operate without multi-agent collaboration, self-consistency checks or external correction loops.
2.3.1 Model Configurations
Pretrained Models: These configurations evaluate the model’s intrinsic few-shot ICL capability using QP-CoT without parameter updates.
Teacher Model: GPT-4o () to establish the upper bound.
Student Model: Qwen3-4B-Instruct-2507 to establish the lower bound.
Tuned Student Models: These configurations evaluate the Student Model after tuning on specific signals.
FN-Gold The Student Model was finetuned using the Gold SQL query () on the BIRD training dataset, with a basic system instruction that directly maps natural language to SQL [pourreza2023dinsql].
ReasonSQL (unstructured KD): The Student Model was finetuned on the complete Teacher sequence , where is the free-form CoT rationale [rossiello2025rationalization]. This configuration serves as the established baseline for evaluating the efficacy of standard unstructured reasoning distillation.
Struct-SQL (Structured KD): The Student Model was finetuned on the entire Teacher sequence , where is the formal query plan CoT trace. This trace represents the structured logical blueprint designed for the core hypothesis.
During inference, all models use the QP-CoT prompt structure except for ReasonSQL, which employs the unstructured CoT prompt.
2.3.2 Distillation Dataset Construction
The KD datasets were constructed using an active generation and filtering methodology to maximize data quality and the diversity of query complexity. To initiate the process, the databases in the corpus are partitioned into a 75% "in-domain" (ID) database pool and a 25% "out-of-domain" (OOD) database pool, based on unique database identifiers. This partitioning ensures that the ID pool serves as the source for all training data and the in-domain validation set, while the OOD pool is used exclusively for the out-of-domain validation set, guaranteeing a robust measure of the trained model’s generalization capabilities.
We then applied stratified, success-based sampling. The ID corpus is iterated using predefined SQL structure and syntax categories to achieve target distributions for the training set (1,000 samples) and the in-domain validation set (150 samples). These categories, as shown in Table 1, are differentiated by the complexity of the SQL structure, including single-table queries, queries with subqueries, queries with joins/set operations and queries that combine both joins/set operations and subqueries. For each candidate, inference was performed using GPT-4o. A sample is admitted only if the generated SQL is both syntactically valid and yields the correct result upon execution. New samples are generated until 1,000 successful training samples and 150 ID validation samples are obtained. The OOD validation set (150 samples) is constructed using the same sampling and validation process, but applied exclusively to the segregated OOD database pool to ensure that no schema overlaps. For both Struct-SQL and ReasonSQL, we used the same data generation pipeline to ensure a controlled, methodologically fair comparison. The datasets differ only in the format of the supervisory signal: unstructured CoT traces for ReasonSQL versus structured QP-CoT for Struct-SQL.
| Complexity Category | Training Count | ID Val Count | OOD Val Count |
| Single Table Queries | 295 (29.50%) | 37 (24.67%) | 46 (30.67%) |
| Subquery (no join or set operations) | 229 (22.90%) | 39 (26.00%) | 31 (20.67%) |
| With JOINs / Set Ops (no Subquery) | 398 (39.80%) | 57 (38.00%) | 60 (40.00%) |
| JOINs / Set Ops and Subquery | 78 (7.80%) | 17 (11.33%) | 13 (8.67%) |
| Total Successful Samples | 1000 | 150 | 150 |
2.4 Implementation and Evaluation Details
2.4.1 Post-Training Details
FN-Gold, ReasonSQL, and Struct-SQL were finetuned using Parameter-Efficient Fine-Tuning (PEFT) using Quantized Low-Rank Adaptation (QLoRA) [dettmers2023qlora] to improve training efficiency. PEFT was chosen for its proven effectiveness in adapting SLMs to new domains with small training datasets. Furthermore, PEFT has been shown to yield more stable models and mitigate the risk of catastrophic forgetting of foundational knowledge. QLoRA allows recovery of the full 16-bit finetuning task performance even when the base model is loaded at 4-bit precision [han2024parameterefficient]. The KD approach keeps the original model parameters fixed and only updates a small set of LoRA adapter parameters . We followed the QLoRA methodology [dettmers2023qlora] and applied adapters to all linear layers of the Transformer architecture. Based on preliminary testing on a subset of the development data, we selected and an alpha scaling of to ensure robust adaptation. Optimization was performed using AdamW with a learning rate of , a maximum input length of tokens, and a generation limit of tokens. We utilized a batch size of 6 for ReasonSQL and Struct-SQL, compared to 15 for the FN-Gold baseline on an NVIDIA H200 GPU. All models minimized completion loss. To balance in-domain accuracy with out-of-domain generalization, we used an early-stopping strategy (patience=8) that monitored the aggregated validation loss across both the ID and OOD validation sets.
2.4.2 Evaluation Metrics
The primary metric is Execution Accuracy (EX), which measures the percentage of generated SQL queries that execute without error and return the same result set as the ground-truth SQL query on the BIRD mini-dev set. A detailed analysis of model failure modes reveals why and where certain models perform best. To facilitate analysis, we categorize failures into three distinct types as defined in Table 2. This classification establishes a severity hierarchy for Text-to-SQL tasks: starting with the most severe Generation Failure (GEN), where the model produces no recognizable SQL output; followed by Syntactic Failure (SYN), indicating an unexecutable query due to grammar or schema errors; and finally, Logical Failure (SEM), representing a query that is syntactically correct but semantically inaccurate. Recognizing this progression is beneficial for discerning the specific challenges and improvements associated with each model.
| Error Type | Subcategory | Description |
| Semantic Errors (SEM) | Incorrect Columns/Rows | Wrong columns/row count |
| Value Mismatch | Incorrect data values | |
| Empty Output | Empty result set | |
| Syntactic Errors (SYN) | No Such Column/Table | Non-existent table/column |
| Keyword Issue | Incorrect or misplaced SQL keyword | |
| Syntax/Clause Order | Incorrect clause order, missing parentheses, etc | |
| Generation Error (GEN) | N/A | No SQL output |
3 Results
3.1 Overall Performance
The execution accuracy of Struct-SQL, compared to all baselines on the BIRD mini-dev dataset, provides strong support for the central hypothesis. As summarized in Table 3, distilling structured reasoning results in significantly better performance compared to unstructured distillation, ReasonSQL and traditional finetuning FN-Gold. The native Student Model’s performance (17.0%) illustrates the Performance Trade-Off in the Adoption Trilemma, indicating that an SLM is insufficient for production deployment without intervention. Struct-SQL achieved an -point absolute improvement from % to % over the ReasonSQL baseline. This improvement enables the Student Model to cover 84% of the Teacher Model’s execution accuracy.
| Overall | EX by Difficulty | |||||
| Model | Prompt | EX | Avg. Tokens | Simple | Moderate | Challenging |
| Teacher (GPT-4o) | QP-CoT | 53.60% | 68.24% | 52.00% | 36.27% | |
| Student Model | QP-CoT | 17.00% | 34.45% | 9.60% | 9.80% | |
| Tuned Student Models | ||||||
| FN-Gold | QP-CoT | 34.30% | 45.94% | 32.80% | 20.58% | |
| ReasonSQL | CoT | 36.90% | 49.32% | 33.20% | 27.45% | |
| Struct-SQL (Ours) | QP-CoT | 45.00% | 65.54% | 40.4% | 25.49% | |
3.2 Analysis of Model Failures
Figure 3 presents hierarchical pie charts. The inner ring displays top-level results that depict success versus categorized error types, while the outer rings further specify subcategories of failures, as defined in Table 2. For successful queries, the outer ring categorizes results by the difficulty level of the SQL queries in the BIRD dataset.
As visually summarized in the hierarchical pie charts, a key dichotomy emerges between the Teacher and other baselines, primarily distinguished by their dominant error types. The Teacher Model (), shown in Figure 3(a), achieved the highest success rate. Its failures were predominantly semantic errors (30%), indicating challenges with logical complexity. The Student Model (), shown in Figure 3(b), had a higher number of syntactic errors, including schema hallucination, underscoring a fundamental inability to adhere to the database schema before post-training.
The standard finetuning baseline (FN-Gold) reflects the performance achieved by training solely on the gold SQL(), without incorporating intermediate reasoning. As shown in Figure 3(c), this approach focused its gains almost entirely on reducing semantic errors (from 53.8% to 33.0%, a -point drop) of the Student Model , but failed to address the syntactic bottleneck: overall syntactic errors remained nearly unchanged ( to ). These results confirm that training on only the final query output is insufficient for learning strict SQL grammar.
The unstructured distillation (ReasonSQL) baseline evaluates distillation through an unstructured reasoning trace (). As shown in Figure 3(d), this approach achieved an execution accuracy of , higher than that of and FN-Gold, indicating that unstructured CoT reasoning provides a better distillation signal than both the naive model and FN-Gold. ReasonSQL significantly reduced the rate of generation errors compared to the naive Student Model , suggesting that unstructured distillation enhanced the stability of the model. Furthermore, it reduced semantic errors from 53.8% to 39.8% (a -point reduction) and reduced syntactic errors from to (a -point reduction), primarily by improving basic schema linking and eliminating ‘No Such Table’ errors. These findings align with previous research demonstrating that CoT-based distillation improves accuracy [hsieh2023cotdistilling] and further clarify its specific advantages.
As established above, unstructured distillation ReasonSQL outperformed both the naive Student Model and standard finetuning FN-Gold. Struct-SQL, as shown in Figure 3(e), outperformed ReasonSQL. By distilling the reasoning signal into a structured blueprint, Struct-SQL increased execution accuracy from to (an -point improvement) compared to ReasonSQL. This improvement appeared across the error severity hierarchy. Struct-SQL nearly eliminated generation errors, reducing the rate from to (a -point reduction). For syntactic errors, although ReasonSQL outperformed the naive model, it did not enforce strict schema alignment. In contrast, Struct-SQL reduced the total syntactic error count from in ReasonSQL to (a -point reduction). This reduction included fewer ‘No Such Column’ hallucinations, to (a -point reduction) and the elimination of ‘Keyword Issues’, showing a more precise understanding of SQL syntax. For semantic errors, the structured model showed greater logical reliability, reducing ‘Empty Output’ errors from in ReasonSQL to (a -point reduction). Although this increased rigidity led to a minor trade-off, ReasonSQL retained a slight advantage in ‘Value Mismatches’ ( vs. ), suggesting that the structured plan provides a more robust signal.
3.3 Fine-Grained Performance Analysis and Ablation Study
To investigate the precise source of performance of Struct-SQL, we move beyond aggregate metrics to examine the detailed differences between models.
Performance by SQL Construct: Assessing model performance across the spectrum of query formulations can be insightful; therefore, we analyzed execution accuracy for key SQL constructs, as shown in Figure 4(a). The comparative analysis indicates that Struct-SQL’s performance gains are most pronounced in tasks that require aggregation and explicit structural decomposition. This is evidenced by its strong results on specific SQL operations. For queries requiring a GROUP BY clause, the Struct-SQL model ( EX) outperformed both FN-Gold ( EX) and ReasonSQL ( EX), demonstrating that explicitly distilling the query execution plan effectively trains the model on the necessary aggregation formulation. An advantage was also observed for queries requiring an ORDER BY, where Struct-SQL showed a -point absolute improvement over ReasonSQL. However, the comparative analysis revealed a notable exception: for queries that require JOIN operations, ReasonSQL achieved the highest execution accuracy. In contrast, Struct-SQL’s score ( EX) was lower, suggesting that the unstructured CoT trace provided a unique advantage for learning multi-entity linking. This area remains a core challenge for all models, as confirmed by the Teacher Model’s performance ( EX). Furthermore, all baselines, including Struct-SQL, shared a weakness in handling queries that require Set Operations.
| Overall | EX by Difficulty | |||
| Model | EX (%) | Simple | Moderate | Challenging |
| Student Model | 7.22 | 11.49 | 6.00 | 3.92 |
| ReasonSQL (CoT) | 25.10 | 40.54 | 20.08 | 12.75 |
| Struct-SQL (Ours) | 29.31 | 49.32 | 23.20 | 14.71 |
Evaluation of Knowledge Transfer: Gains vs. Losses: A Gains vs. Losses analysis was conducted to assess the quality of the distilled knowledge (Figure 4(b)). This approach compares queries where the Teacher was correct, but the Student Model failed (Losses, red segments) with cases where the Student Model was correct, but the Teacher Model was incorrect (Gains, green segments), representing true generalization. The results indicate that Struct-SQL demonstrates the highest net performance relative to the Teacher Model by exhibiting the fewest losses and the most gains among all student models. This confirms that the structured reasoning signal leads to better fidelity in knowledge transfer.
Performance State Transitions: Figure 4(c) illustrates the progression of queries from the baseline Student Model to the distilled Struct-SQL, highlighting the systematic conversion of error states into successful or unsuccessful output. Struct-SQL retained more than 81% of the initial successes and enhanced stability by reducing generation errors of the Student Model from 4.0% to 0.4%. The structured framework also corrected 41.3% of semantic errors (SEM SUCC), confirming that the formal query plan addresses logical flaws. Additionally, over two-thirds of syntactic failures were upgraded: 29% were corrected (SYN SUCC) and 41% shifted to semantic errors (SYN SEM).
Generalization on Mistral-7B: To validate the transferability of our framework to other architectures, we replicated our experiments on Mistral-7B-Instruct-v3.0. This model was specifically selected for its low zero-shot accuracy of 4.4% EX on BIRD mini-dev. As shown in Table 4, Struct-SQL shows generalizability, achieving a performance advantage (29.3% EX) over the Student Model (7.2% EX) and the ReasonSQL baseline (25.1% EX), confirming that the structured query plan offers a robust supervision signal independent of the base model.
Computational Efficiency: On a single H200 GPU, using Qwen3-4B-Instruct-2507 as the Student Model, Struct-SQL converged using an early stopping strategy with a patience of 8 steps and a threshold of 0.001 in 2.24 epochs, requiring only 29.15 minutes. This training efficiency is comparable to that of the unstructured ReasonSQL, which required 25.24 minutes across 6.4 epochs with the same batch size, sample count and early stopping strategy. This 1,000-sample distillation strategy is also computationally tractable compared to traditional finetuning on larger datasets. For context, the FN-Gold baseline, trained on the entire 9,000+ sample BIRD dataset, required 110.57 minutes (4.33 epochs) to converge. This efficiency underscores the practical deployability of our structured distillation method in resource-constrained environments.
Official Test Evaluation: Verified by the BIRD benchmark team, our model achieves 60.42% EX333BIRD public leaderboard listing is deferred to maintain double-blind anonymity., ranking first among B models (as of December 18, 2025). This result is achieved using strict single-model inference with greedy decoding and no self-consistency.
4 Related Work
4.1 Text-to-SQL with In-Context Learning (ICL)
LLMs fundamentally redefined the field of Text-to-SQL [deng2022recent], achieving substantial performance on challenging cross-domain benchmarks such as Spider [yu2018spider] and BIRD [li2024can]. ICL [dong-etal-2024-survey] has emerged as a dominant research paradigm, enabling LLMs to leverage their pre-training by processing instructions and examples for SQL generation [shi2025asurvey]. Initially, ICL approaches utilized foundational zero-shot and few-shot paradigms [dong2023zero, rajkumar2022evaluating, liu2023comprehensive, wang-etal-2025-mac]. However, generating complex SQL queries that involve nested logic and multitable joins was found to require explicit external guidance [pourreza2023dinsql]. As a result, the field pivoted toward decomposition, a multi-step reasoning tactic designed to improve accuracy and reliability [pourreza2023dinsql, zhang2023act, gao2024dail, Li2024pet, liu2023divide]. As a form of ICL-based decomposition, basic CoT helped LLMs break down the task, follow a logical workflow and improve execution accuracy [liu2023divide]. This progression led to sophisticated variants such as DIN-SQL [pourreza2023dinsql], which decomposes the task into sequential stages (e.g., schema linking, query classification). Additional multi-stage architectural interventions have also been explored [zhang2023act, Li2024pet]. Crucially, these approaches often require multiple LLM calls to intervene, correct, or gather intermediate data, leading to high latency and cost. To mitigate this, some approaches focused on optimizing efficiency by creating single-pass Structural CoT approaches [pourreza2024chase, liu2023divide]. These methods instruct the LLM to follow a formalized logical blueprint to gain the necessary understanding of the schema and natural language before generating the SQL. Specialized methods, including QP-CoT and QDecompose, excel here, as they have been shown to produce better results than non-structured CoT [pourreza2024chase, gao2024dail]. For example, QP-CoT provides few-shot query plan examples that require the model to first generate a detailed query plan before generating SQL code [pourreza2024chase]. Although single-pass ICL performs well with large LLMs, its gains are limited when applied to SLMs [maamari2024death] or require expensive finetuning [pourreza2024dts, liu2025uncovering, gorti-etal-2025-msc, zhai2025excot]. This work addresses this identified gap by introducing Struct-SQL, a novel KD method that enables single-pass ICL for SLMs.
4.2 KD for Reasoning Transfer
KD has evolved from a model compression technique to a method for complex knowledge transfer, called Skill Distillation, often used to transfer reasoning knowledge from one model to others or to facilitate self-improvement [hinton2014distilling, shi2025asurvey]. Due to the significant resource demands of LLMs, recent research has focused on efficiently transferring its reasoning capabilities from a teacher LLM to a student SLM. This transfer is frequently grounded in ICL principles, demonstrating that learning from a few-shot demonstrations can be successfully distilled into the SLMs’ parameters [snell2022learning]. The primary established approach to enable reasoning in SLMs is the Finetune-CoT paradigm [li2024explanations], which uses unstructured CoT explanations of the teacher as a supervision signal. This distillation approach has been demonstrated to be effective with multi-step mathematical reasoning [yao2023specializing]. However, relying on unstructured CoT often leads SLMs to learn spurious correlations rather than deep causal features, thereby reducing robustness on complex data [li2023mixed, shridhar2023distilling]. Consequently, recent research increasingly favors methods that enforce structural decomposition and explicit causality. These approaches include creating constrained distillation pipelines, with the aim of removing ambiguous input context through advanced structural formats. For example, SocraticCoT provides an explicit structural format by guiding the model through reasoning using defined subquestion-solution pairs [shridhar2023distilling]. Alternatively, Mixed Distillation integrates the CoT with a formal verifiable Program-of-Thought reasoning to provide a less ambiguous supervision signal [li2023mixed]. In the Text-to-SQL domain, structured KD has seen limited uptake. Previous studies, while validating the effectiveness of distillation in this area, have generally relied on less structured signals, such as schema-based finetuning to improve schema linking [hong2024next], the inclusion of enterprise-specific custom examples [hoang-etal-2025-distill], and unstructured CoT reasoning transfer [rossiello2025rationalization]. The proposed Struct-SQL framework systematically evaluates whether structure-based reasoning transfer, specifically QP-CoT, improves the distillation for Text-to-SQL.
5 Conclusion and Future Research
This work presents Struct-SQL, a KD framework that transfers structured reasoning (QP-CoT) from a Teacher LLM to a smaller student model for the Text-to-SQL task. The central hypothesis is that a formal, structured reasoning signal provides a superior, less ambiguous teaching blueprint than unstructured reasoning. The experiments confirm the hypothesis, demonstrating that distilling structured reasoning is a better teaching method. Detailed error analysis indicates that these improvements are primarily due to a significant reduction in syntactic errors. An additional ablation study demonstrated generalization to a different base model. This increased precision, however, comes with a trade-off: the mandatory generation of the intermediate query plan increases the inference token overhead and relies on a fixed reasoning template. Additionally, the Student’s performance remains bound by the Teacher’s fidelity, particularly with respect to complex join and set operations. Future research should therefore explore integrating or combining this structured reasoning with other established strategies to determine the optimal signal for distillation efficiency. Given the demonstrated effectiveness of structured signals in the refinement of syntactic efficiency and logical reasoning, further research is warranted to validate the effect of this structured distillation approach on other complex reasoning tasks beyond Text-to-SQL. In summary, the proposed framework facilitates the deployment of high-performing, cost-effective and private models.
6 Acknowledgments
We thank Khalid Eidoo, Abdul Hamid Dabboussi, María Rodríguez-Liñán and Ting Fung Lam for their insightful technical discussions and valuable feedback throughout the project and during the review of this manuscript.

### v2

@@
 To evaluate this hypothesis, we propose Struct-SQL, a novel Knowledge Distillation (KD) framework that trains an SLM to emulate a powerful large LLM.
-Consequently, we adopt a query execution plan as a formal blueprint to derive this structured reasoning.
-Our SLM, distilled with structured CoT, achieves an absolute improvement of 8.1% over an unstructured CoT distillation baseline.
+To implement this approach, we adopt the query execution plan as a formal blueprint to derive structured reasoning.
+Our SLM, distilled with a structured CoT, achieves an absolute improvement of 8.1% over an unstructured CoT distillation baseline.
 A detailed error analysis reveals that a key factor in this gain is a marked reduction in syntactic errors.
@@
 1 Introduction
-Text-to-SQL (NL2SQL or text2sql) has the potential to democratize data access [fei2014constructing].
-The field has seen substantial performance advancements driven by the advent of Large Language Models (LLMs) [shi2025asurvey].
-These models enhance the capabilities of natural language interfaces for databases by automatically translating natural-language user questions into SQL queries [hong2024next].
+Text-to-SQL (NL2SQL or text2sql) has the potential to democratize data access [14].
+The field has seen substantial performance advancements driven by the advent of Large Language Models (LLMs) [28].
+These models enhance the capabilities of natural language interfaces for databases by automatically translating natural-language user questions into SQL queries [11].
 Nevertheless, widespread adoption in enterprises remains challenging due to a difficult trade-off among three interdependent factors: cost, security, and performance.
 This challenge can be understood as an Adoption Trilemma.
-Cost: High-performing models typically require significant computational resources, leading to high operational costs, whether using proprietary APIs or privately hosted LLMs [shi2025asurvey].
-Security: Relying on external APIs raises prominent security concerns, as transmitting potentially sensitive database schemas and sample records to third-party providers is often unacceptable in enterprise settings [hoang-etal-2025-distill].
-Performance: Selecting open-source models for local, private deployment to address cost and security issues often leads to the use of Small Language Models (SLMs), which typically lack adequate zero-shot accuracy for complex real-world queries [li2024can].
+Cost: High-performing models typically require significant computational resources, leading to high operational costs, whether using proprietary APIs or privately hosted LLMs [28].
+Security: Relying on external APIs raises serious security concerns, as transmitting potentially sensitive database schemas and sample records to third-party providers is often unacceptable in enterprise settings [10].
+Performance: Selecting open-source models for local, private deployment to address cost and security issues often leads to the use of Small Language Models (SLMs), which typically lack adequate zero-shot accuracy for complex real-world queries [15].
 To identify the limitations of current approaches, we examine the impact of recent advances in reasoning-driven prompting on Text-to-SQL performance.
-A notable portion of the recent performance gain in Text-to-SQL using LLMs can be attributed to in-context learning (ICL) [dong-etal-2024-survey].
-ICL-based methods, particularly those that use decomposition and multi-step reasoning, have demonstrated substantial gains in performance on the Text-to-SQL task [hong2024next].
-A prominent ICL technique for this is Chain-of-Thought (CoT), which encourages models to think step-by-step [wei2022chain, liu2025uncovering].
+A notable portion of the recent performance gains in Text-to-SQL using LLMs can be attributed to in-context learning (ICL) [3].
+ICL-based methods, particularly those that use decomposition and multi-step reasoning, have demonstrated substantial gains in performance on the Text-to-SQL task [11].
+A prominent ICL technique for this is Chain-of-Thought (CoT), which encourages models to think step-by-step [32, 19].
 Such decomposition strategies are beneficial for generating complex SQL queries.
-For instance, DAIL-SQL [gao2024dail] enhances CoT through context-aware examples, while DIN-SQL [pourreza2023dinsql] and Divide-and-Conquer CoT (DC-CoT) [pourreza2024chase] improve accuracy by breaking questions into intermediate sub-queries.
-Building on this logical foundation, the Query Plan CoT (QP-CoT) guides the model through steps that mirror a database execution plan, simulating the logical process by which databases execute queries.
-By following this structured path, QP-CoT ensures that the SQL generation path aligns with the inherent logic of query execution [pourreza2024chase].
-Although these reasoning techniques achieve considerable success, their effectiveness is observed almost exclusively with large LLMs.
-This reliance on large LLMs constitutes a significant limitation: these methods depend on the very models that exacerbate the cost and security challenges, thereby limiting their applicability in resource-constrained enterprise settings and contributing to the ‘Adoption Trilemma’.
+For instance, DAIL-SQL [6] enhances CoT through context-aware examples, while DIN-SQL [23] and Divide-and-Conquer CoT (DC-CoT) [22] improve accuracy by breaking questions into intermediate sub-queries.
+Building on this logical foundation, the Query Plan CoT (QP-CoT) guides the model through the steps that mirror a database execution plan, simulating the logical process by which databases execute queries.
+By following this structured path, QP-CoT ensures that the SQL generation path aligns with the inherent logic of query execution [22].
+Although these reasoning techniques achieve considerable success, their effectiveness is observed almost exclusively in the large LLMs.
+This reliance on large LLMs constitutes a significant limitation: these methods depend on the very models that exacerbate the cost and security challenges, thereby limiting their applicability to resource-constrained enterprise settings and contributing to the ‘Adoption Trilemma’.
 The performance limitations of SLMs, suitable for private deployment, are particularly pronounced in the Text-to-SQL domain.
 Standard benchmarks highlight a significant gap between large models and smaller open-source alternatives.
-For example, on the BIRD mini-dev benchmark, LLMs such as GPT-4o and Claude (150B parameters) achieve execution accuracies between 30% and 45%, whereas widely used SLMs such as Mistral, Mixtral and Qwen2.5 Coder (7B parameters) achieve only 4% to 12%111Performance on BIRD mini-dev, see https://github.com/bird-bench/mini_dev.
+For example, on the BIRD mini-dev benchmark, LLMs such as GPT-4o and Claude 3.7 Sonnet (150B parameters) achieve execution accuracies between 30% and 45% using standard prompting, whereas widely used SLMs such as Mistral, Mixtral and Qwen2.5 Coder (7B parameters) achieve only 4% to 12%111Performance on BIRD mini-dev, see https://github.com/bird-bench/mini_dev.
 This paper reveals that this severe performance degradation in SLMs persists even when advanced reasoning prompts are employed, despite their effectiveness in larger LLMs.
-This finding is consistent with recent literature, suggesting that large LLMs can adhere to the schema as long as the schema fits within their context window, a capability not observed in SLMs [maamari2024death].
+This finding is consistent with recent literature, suggesting that large LLMs can adhere to the schema as long as the schema fits within their context window, a capability not observed in SLMs [21].
 In our experiments, the SLM (Qwen3-4B-Instruct-2507) prompted with QP-CoT fails primarily due to inadequate schema adherence, with a pronounced tendency toward schema hallucination, often generating non-existent tables or columns (see Figure 3(b)).
@@
 This breakdown in structured reasoning motivates an investigation into whether the structure of reasoning itself can be effectively transferred from large LLMs to SLMs.
-To bridge this performance gap, Knowledge Distillation (KD) [hinton2014distilling], which transfers reasoning ability from a capable teacher model, is a key strategy [shi2025asurvey].
+To bridge this performance gap, Knowledge Distillation (KD) [9], which transfers reasoning ability from a capable teacher model, is a key strategy [28].
 KD is a model compression technique in which a smaller "student" model is trained to mimic the behavior of a larger, pretrained "teacher" model.
-Beyond compression, KD enables the transfer of complex reasoning [li2024explanations] and logical skills [hsieh2023cotdistilling, yao2023specializing].
+Beyond compression, KD enables the transfer of complex reasoning [16] and logical skills [12, 5].
 Through KD, our aim is to build SLMs for Text-to-SQL that deliver accuracy comparable to larger LLMs while meeting enterprise cost and security requirements through private deployment.
-This work argues that KD is a promising approach in addressing this trilemma, but its effectiveness depends on the type of reasoning being distilled.
+This work argues that KD is a promising approach to addressing this trilemma, but its effectiveness depends on the type of reasoning being distilled.
 Given that the core bottleneck for SLMs is replicating the teacher’s complex reasoning, the central question for KD is identifying the optimal form of the reasoning signal to distill.
-Building on the success of CoT prompting [wei2022chain], one approach is to distill the teacher’s natural language reasoning traces [li2024explanations].
+Building on the success of CoT prompting [32], one approach is to distill the teacher’s natural language reasoning traces [16].
 This approach is exemplified by ReasonSQL, where the student model is trained on the teacher’s intermediate CoT steps in addition to the final SQL query.
-This unstructured approach provides a better knowledge signal than finetuning on SQL queries alone and has been shown to improve model accuracy [rossiello2025rationalization].
+This unstructured approach provides a better knowledge signal than finetuning on SQL queries alone and has been shown to improve model accuracy [27].
 However, we argue that the structure of the reasoning signal itself is critical for effective distillation.
@@
 Figure 1 shows the KD workflow, contrasting the proposed structured reasoning approach against the standard unstructured CoT method.
-Within this framework, a state-of-the-art (SOTA) Teacher Model is used to generate a QP-CoT trace, which formally decomposes the query into a logical execution plan [pourreza2024chase].
+Within this framework, a state-of-the-art (SOTA) Teacher Model is used to generate a QP-CoT trace, which formally decomposes the query into a logical execution plan [22].
 This structured plan, together with the generated SQL query, constitutes the supervisory signal.
@@
 During inference, this structured reasoning is retained: the student model is given the QP-CoT prompt to autonomously generate the query plan before synthesizing the final SQL query.
-We perform an extensive comparative analysis on the mini-dev BIRD benchmark [li2024can], evaluating our Struct-SQL framework against key baselines. Our contributions are as follows.
+To validate our hypothesis, we perform an extensive comparative analysis on the mini-dev BIRD benchmark [15], evaluating our Struct-SQL framework against key baselines.
+Our main contributions are as follows.
 We are the first to systematically evaluate the impact of KD using a structured reasoning signal for Text-to-SQL.
-We provide a comprehensive error analysis, showing that the improvement in Struct-SQL results from a marked reduction in syntactic errors (e.g., schema hallucination), demonstrating that the structured signal provides a clearer curriculum.
+We provide a comprehensive error analysis, showing that the improvement in Struct-SQL results from a marked reduction in the syntactic errors (e.g., schema hallucination), demonstrating that the structured signal provides a clearer curriculum.
 We validate the generalization of our framework through an ablation study on two different SLMs.
-We release the KD dataset of 1,300 structured reasoning traces, the Struct-SQL model, prompts and the code 222https://github.com/craterlabs/Struct-SQL-Distillation/.
+We release the code (github.com/craterlabs/Struct-SQL-Distillation) and, the model and the dataset (huggingface.co/collections/craterlabs/struct-sql) to facilitate reproducible research.
 Following this introduction, Section 2 details the methodology. Section 3 presents the results, which are contextualized by the related work in Section 4, and the paper concludes in Section 6.
@@
 A query execution plan defines the precise sequence of steps that a database follows to access and manipulate data, often generated using an EXPLAIN command.
-We adopt this ICL prompt strategy from a recent work on a Query Execution Plan-based prompting strategy, QP-CoT [pourreza2024chase].
+We adopt this ICL prompt strategy from a recent work on a Query Execution Plan-based prompting strategy, QP-CoT [22].
 Instead of a free-form explanation, the Teacher Model is prompted to generate a query plan along with the final SQL query.
 As shown in Figure 2, this plan decomposes the query into a sequential execution flow that explicitly performs selections, filters, joins, and groupings through step-by-step table scanning and data manipulation, providing a systematic signal for the student model.
-Although the query plan has been explored as a prompting technique [pourreza2024chase], the key contribution here is to demonstrate the effectiveness of this structured reasoning as the primary teaching signal within a KD framework.
+Although the query plan has been explored as a prompting technique [22], the key contribution here is to demonstrate the effectiveness of this structured reasoning as the primary teaching signal within a KD framework.
 2.3 Experimental Setup
-The experiments use the SQLite-based BIRD benchmark for the data [li2024can].
+The experiments use the SQLite-based BIRD benchmark for the data [15].
 The BIRD training dataset is used for model training, while the BIRD mini-dev dataset serves for evaluation.
@@
 The Student Model () is Qwen3-4B-Instruct-2507 (Alibaba Cloud).
-It was chosen for its strong performance-to-size ratio, which enables low-latency, private deployment and thus addresses the Adoption Trilemma [yang2025qwen3].
+It was chosen for its strong performance-to-size ratio, which enables low-latency, private deployment and thus addresses the Adoption Trilemma [33].
 All models utilize a single-pass inference mechanism and operate without multi-agent collaboration, self-consistency checks or external correction loops.
+For a fair comparison with baselines, no external reasoning signals were provided at test time.
+During inference, the student model autonomously generates the Query Execution Plan based on the input prompt before generating the final SQL.
 2.3.1 Model Configurations
@@
 Student Model: Qwen3-4B-Instruct-2507 to establish the lower bound.
-Tuned Student Models: These configurations evaluate the Student Model after tuning on specific signals.
-FN-Gold The Student Model was finetuned using the Gold SQL query () on the BIRD training dataset, with a basic system instruction that directly maps natural language to SQL [pourreza2023dinsql].
-ReasonSQL (unstructured KD): The Student Model was finetuned on the complete Teacher sequence , where is the free-form CoT rationale [rossiello2025rationalization].
+Tuned Student Models: These configurations evaluate the Student Model after tuning it on specific signals.
+FN-Gold The Student Model was finetuned using the Gold SQL query () on the BIRD training dataset, with a basic system instruction that directly maps natural language to SQL [23].
+ReasonSQL (unstructured KD): The Student Model was finetuned on the complete Teacher sequence , where is the free-form CoT rationale [27].
 This configuration serves as the established baseline for evaluating the efficacy of standard unstructured reasoning distillation.
@@
 This partitioning ensures that the ID pool serves as the source for all training data and the in-domain validation set, while the OOD pool is used exclusively for the out-of-domain validation set, guaranteeing a robust measure of the trained model’s generalization capabilities.
-We then applied stratified, success-based sampling.
+We then applied stratified, success-based sampling, where samples are grouped by SQL complexity categories and only admitted if the Teacher Model generates syntactically valid and execution-correct SQL.
 The ID corpus is iterated using predefined SQL structure and syntax categories to achieve target distributions for the training set (1,000 samples) and the in-domain validation set (150 samples).
@@
 2.4.1 Post-Training Details
-FN-Gold, ReasonSQL, and Struct-SQL were finetuned using Parameter-Efficient Fine-Tuning (PEFT) using Quantized Low-Rank Adaptation (QLoRA) [dettmers2023qlora] to improve training efficiency.
+FN-Gold, ReasonSQL, and Struct-SQL were finetuned using Parameter-Efficient Fine-Tuning (PEFT) using Quantized Low-Rank Adaptation (QLoRA) [2] to improve training efficiency.
 PEFT was chosen for its proven effectiveness in adapting SLMs to new domains with small training datasets.
 Furthermore, PEFT has been shown to yield more stable models and mitigate the risk of catastrophic forgetting of foundational knowledge.
-QLoRA allows recovery of the full 16-bit finetuning task performance even when the base model is loaded at 4-bit precision [han2024parameterefficient].
+QLoRA allows recovery of near full 16-bit finetuning task performance even when the base model is loaded at 4-bit precision [8].
 The KD approach keeps the original model parameters fixed and only updates a small set of LoRA adapter parameters .
-We followed the QLoRA methodology [dettmers2023qlora] and applied adapters to all linear layers of the Transformer architecture.
+We followed the QLoRA methodology [2] and applied adapters to all linear layers of the Transformer architecture.
 Based on preliminary testing on a subset of the development data, we selected and an alpha scaling of to ensure robust adaptation.
@@
 To facilitate analysis, we categorize failures into three distinct types as defined in Table 2.
-This classification establishes a severity hierarchy for Text-to-SQL tasks: starting with the most severe Generation Failure (GEN), where the model produces no recognizable SQL output; followed by Syntactic Failure (SYN), indicating an unexecutable query due to grammar or schema errors; and finally, Logical Failure (SEM), representing a query that is syntactically correct but semantically inaccurate.
+This classification establishes a severity hierarchy for Text-to-SQL tasks: starting with the most severe Generation Failure (GEN), where the model produces no recognizable SQL output; followed by Syntactic Failure (SYN), which indicates an unexecutable query due to grammar or schema errors; and finally, Logical Failure (SEM), representing a query that is syntactically correct but semantically inaccurate.
 Recognizing this progression is beneficial for discerning the specific challenges and improvements associated with each model.
 | Error Type | Subcategory | Description |
-| Semantic Errors (SEM) | Incorrect Columns/Rows | Wrong columns/row count |
+| Generation (GEN) | — | No SQL output |
+| Syntactic (SYN) | No Such Column | Non-existent column reference |
+| No Such Table | Non-existent table reference | |
+| Keyword Issue | Incorrect or misplaced SQL keyword | |
+| Syntax/Clause Order | Incorrect clause order, missing parentheses | |
+| Other | Unclassified syntactic errors | |
+| Semantic (SEM) | Column Mismatch | Incorrect number of columns |
+| Row Mismatch | Incorrect number of row | |
+| Row & Column Mismatch | Incorrect number of columns and rows | |
 | Value Mismatch | Incorrect data values | |
-| Empty Output | Empty result set | |
-| Syntactic Errors (SYN) | No Such Column/Table | Non-existent table/column |
-| Keyword Issue | Incorrect or misplaced SQL keyword | |
-| Syntax/Clause Order | Incorrect clause order, missing parentheses, etc | |
-| Generation Error (GEN) | N/A | No SQL output |
+| Empty Output | Empty result set |
 3 Results
@@
 | ReasonSQL | CoT | 36.90% | 49.32% | 33.20% | 27.45% | |
+| ReasonSQL (mismatched)222Distilled using CoT, evaluated with QP-CoT. | QP-CoT | 29.20% | 46.62% | 24.80% | 14.71% | |
 | Struct-SQL (Ours) | QP-CoT | 45.00% | 65.54% | 40.4% | 25.49% | |
+To isolate whether Struct-SQL’s performance gains stem from the structured QP-CoT prompt format at inference time or from the structured supervisory signal during training, we conducted an additional ablation study.
+We evaluated the ReasonSQL model—trained on unstructured CoT traces—using the QP-CoT prompt at inference, denoted as ReasonSQL (mismatched) in Table 3.
+This configuration achieved only 29.2% execution accuracy, a substantial drop from the 36.9% achieved when ReasonSQL uses its native CoT prompt.
+This 7.7-point degradation reveals a critical prompt-training mismatch: a model trained on unstructured reasoning cannot effectively leverage a structured prompt format, even when the prompt includes few-shot examples demonstrating how to generate the structured query plan.
+This suggests that in SLMs, ICL alone is insufficient to bridge the gap; rather, the model must be explicitly trained on structured reasoning to internalize the required logical decomposition patterns.
+The strength of the training signal is further evidenced by Struct-SQL, which achieves 45.0% EX using the identical QP-CoT prompt.
+This disparity validates that Struct-SQL’s success stems from internalizing the logical decomposition patterns during training, which the prompt alone cannot replicate.
 3.2 Analysis of Model Failures
@@
 Furthermore, it reduced semantic errors from 53.8% to 39.8% (a -point reduction) and reduced syntactic errors from to (a -point reduction), primarily by improving basic schema linking and eliminating ‘No Such Table’ errors.
-These findings align with previous research demonstrating that CoT-based distillation improves accuracy [hsieh2023cotdistilling] and further clarify its specific advantages.
+These findings align with previous research demonstrating that CoT-based distillation improves accuracy [12] and further clarify its specific advantages.
 As established above, unstructured distillation ReasonSQL outperformed both the naive Student Model and standard finetuning FN-Gold.
@@
 To investigate the precise source of performance of Struct-SQL, we move beyond aggregate metrics to examine the detailed differences between models.
-Performance by SQL Construct: Assessing model performance across the spectrum of query formulations can be insightful; therefore, we analyzed execution accuracy for key SQL constructs, as shown in Figure 4(a).
+In this section, we deconstruct the model’s capabilities across three dimensions: its proficiency with complex SQL operations (Figure 4(a)), the fidelity of its alignment with the teacher model (Figure 4(b)), and its ability to systematically rectify the student model’s baseline errors (Figure 4(c)).
+3.3.1 Performance by SQL Construct:
+Assessing model performance across the spectrum of query formulations can be insightful; therefore, we analyzed execution accuracy for key SQL constructs, as shown in Figure 4(a).
 The comparative analysis indicates that Struct-SQL’s performance gains are most pronounced in tasks that require aggregation and explicit structural decomposition.
@@
 | Student Model | 7.22 | 11.49 | 6.00 | 3.92 |
-| ReasonSQL (CoT) | 25.10 | 40.54 | 20.08 | 12.75 |
+| ReasonSQL | 25.10 | 40.54 | 20.08 | 12.75 |
 | Struct-SQL (Ours) | 29.31 | 49.32 | 23.20 | 14.71 |
-Evaluation of Knowledge Transfer: Gains vs.
-Losses: A Gains vs.
+3.3.2 Evaluation of Knowledge Transfer: Gains vs. Losses:
+A Gains vs.
 Losses analysis was conducted to assess the quality of the distilled knowledge (Figure 4(b)).
@@
 This confirms that the structured reasoning signal leads to better fidelity in knowledge transfer.
-Performance State Transitions: Figure 4(c) illustrates the progression of queries from the baseline Student Model to the distilled Struct-SQL, highlighting the systematic conversion of error states into successful or unsuccessful output.
+3.3.3 Performance State Transitions:
+Figure 4(c) illustrates the progression of queries from the baseline Student Model to the distilled Struct-SQL, highlighting the systematic conversion of error states into successful or unsuccessful output.
 Struct-SQL retained more than 81% of the initial successes and enhanced stability by reducing generation errors of the Student Model from 4.0% to 0.4%.
@@
 Additionally, over two-thirds of syntactic failures were upgraded: 29% were corrected (SYN SUCC) and 41% shifted to semantic errors (SYN SEM).
-Generalization on Mistral-7B: To validate the transferability of our framework to other architectures, we replicated our experiments on Mistral-7B-Instruct-v3.0.
-This model was specifically selected for its low zero-shot accuracy of 4.4% EX on BIRD mini-dev.
+3.4 Generalization on Mistral-7B:
+To validate the transferability of our framework to other architectures, we replicated our experiments on Mistral-7B-Instruct-v3.0.
+This model was specifically selected for its low zero-shot accuracy of 7.2% EX on BIRD mini-dev with QP-CoT prompt.
 As shown in Table 4, Struct-SQL shows generalizability, achieving a performance advantage (29.3% EX) over the Student Model (7.2% EX) and the ReasonSQL baseline (25.1% EX), confirming that the structured query plan offers a robust supervision signal independent of the base model.
-Computational Efficiency: On a single H200 GPU, using Qwen3-4B-Instruct-2507 as the Student Model, Struct-SQL converged using an early stopping strategy with a patience of 8 steps and a threshold of 0.001 in 2.24 epochs, requiring only 29.15 minutes.
+3.5 Computational Efficiency:
+On a single H200 GPU, using Qwen3-4B-Instruct-2507 as the Student Model, Struct-SQL converged using an early stopping strategy with a patience of 8 steps and a threshold of 0.001 in 2.24 epochs, requiring only 29.15 minutes.
 This training efficiency is comparable to that of the unstructured ReasonSQL, which required 25.24 minutes across 6.4 epochs with the same batch size, sample count and early stopping strategy.
@@
 This efficiency underscores the practical deployability of our structured distillation method in resource-constrained environments.
-Official Test Evaluation: Verified by the BIRD benchmark team, our model achieves 60.42% EX333BIRD public leaderboard listing is deferred to maintain double-blind anonymity., ranking first among B models (as of December 18, 2025).
+3.6 Official Test Evaluation:
+As shown in Table 4, Struct-SQL obtained 60.42% EX on the non-public BIRD benchmark [15], securing the top position among 4B models (as of January 30, 2026) in single model inference.
 This result is achieved using strict single-model inference with greedy decoding and no self-consistency.
+| Model | Simple (%) | Moderate (%) | Challenging (%) | Overall EX (%) |
+| Struct-SQL(4B) | 69.02 | 54.41 | 43.51 | 60.42 |
 4 Related Work
 4.1 Text-to-SQL with In-Context Learning (ICL)
-LLMs fundamentally redefined the field of Text-to-SQL [deng2022recent], achieving substantial performance on challenging cross-domain benchmarks such as Spider [yu2018spider] and BIRD [li2024can].
-ICL [dong-etal-2024-survey] has emerged as a dominant research paradigm, enabling LLMs to leverage their pre-training by processing instructions and examples for SQL generation [shi2025asurvey].
-Initially, ICL approaches utilized foundational zero-shot and few-shot paradigms [dong2023zero, rajkumar2022evaluating, liu2023comprehensive, wang-etal-2025-mac].
-However, generating complex SQL queries that involve nested logic and multitable joins was found to require explicit external guidance [pourreza2023dinsql].
-As a result, the field pivoted toward decomposition, a multi-step reasoning tactic designed to improve accuracy and reliability [pourreza2023dinsql, zhang2023act, gao2024dail, Li2024pet, liu2023divide].
-As a form of ICL-based decomposition, basic CoT helped LLMs break down the task, follow a logical workflow and improve execution accuracy [liu2023divide].
-This progression led to sophisticated variants such as DIN-SQL [pourreza2023dinsql], which decomposes the task into sequential stages (e.g., schema linking, query classification).
-Additional multi-stage architectural interventions have also been explored [zhang2023act, Li2024pet].
+LLMs fundamentally redefined the field of Text-to-SQL [1], achieving substantial performance on challenging cross-domain benchmarks such as Spider [35] and BIRD [15].
+ICL [3] has emerged as a dominant research paradigm, enabling LLMs to leverage their pre-training by processing instructions and examples for SQL generation [28].
+Initially, ICL approaches utilized foundational zero-shot and few-shot paradigms [4, 26, 18, 31].
+However, generating complex SQL queries that involve nested logic and multitable joins was found to require explicit external guidance [23].
+As a result, the field pivoted toward decomposition, a multi-step reasoning tactic designed to improve accuracy and reliability [23, 37, 6, 17, 20].
+As a form of ICL-based decomposition, basic CoT helped LLMs break down the task, follow a logical workflow, and improve execution accuracy [20].
+This progression led to sophisticated variants such as DIN-SQL [23], which decomposes the task into sequential stages (e.g., schema linking, query classification).
+Additional multi-stage architectural interventions have also been explored [37, 17].
 Crucially, these approaches often require multiple LLM calls to intervene, correct, or gather intermediate data, leading to high latency and cost.
-To mitigate this, some approaches focused on optimizing efficiency by creating single-pass Structural CoT approaches [pourreza2024chase, liu2023divide].
+To mitigate this, some approaches focus on optimizing efficiency by creating single-pass Structural CoT approaches [22, 20].
 These methods instruct the LLM to follow a formalized logical blueprint to gain the necessary understanding of the schema and natural language before generating the SQL.
-Specialized methods, including QP-CoT and QDecompose, excel here, as they have been shown to produce better results than non-structured CoT [pourreza2024chase, gao2024dail].
-For example, QP-CoT provides few-shot query plan examples that require the model to first generate a detailed query plan before generating SQL code [pourreza2024chase].
-Although single-pass ICL performs well with large LLMs, its gains are limited when applied to SLMs [maamari2024death] or require expensive finetuning [pourreza2024dts, liu2025uncovering, gorti-etal-2025-msc, zhai2025excot].
+Specialized methods, including QP-CoT and QDecompose, excel here, as they have been shown to produce better results than non-structured CoT [22, 6].
+For example, QP-CoT provides few-shot query plan examples that require the model to first generate a detailed query plan before generating SQL code [22].
+Although single-pass ICL performs well with large LLMs, its gains are limited when applied to SLMs [21] or require expensive finetuning [24, 19, 7, 36].
 This work addresses this identified gap by introducing Struct-SQL, a novel KD method that enables single-pass ICL for SLMs.
 4.2 KD for Reasoning Transfer
-KD has evolved from a model compression technique to a method for complex knowledge transfer, called Skill Distillation, often used to transfer reasoning knowledge from one model to others or to facilitate self-improvement [hinton2014distilling, shi2025asurvey].
-Due to the significant resource demands of LLMs, recent research has focused on efficiently transferring its reasoning capabilities from a teacher LLM to a student SLM.
-This transfer is frequently grounded in ICL principles, demonstrating that learning from a few-shot demonstrations can be successfully distilled into the SLMs’ parameters [snell2022learning].
-The primary established approach to enable reasoning in SLMs is the Finetune-CoT paradigm [li2024explanations], which uses unstructured CoT explanations of the teacher as a supervision signal.
-This distillation approach has been demonstrated to be effective with multi-step mathematical reasoning [yao2023specializing].
-However, relying on unstructured CoT often leads SLMs to learn spurious correlations rather than deep causal features, thereby reducing robustness on complex data [li2023mixed, shridhar2023distilling].
+KD has evolved from a model compression technique to a method for complex knowledge transfer, called Skill Distillation, which is often used to transfer reasoning knowledge from one model to another or to facilitate self-improvement [9, 28].
+Due to the significant resource demands of LLM, recent research has focused on efficiently transferring their reasoning capabilities from a teacher LLM to a student SLM.
+This transfer is frequently grounded in ICL principles, demonstrating that learning from few-shot demonstrations can be successfully distilled into the SLMs’ parameters [30].
+The primary established approach to enable reasoning in SLMs is the Finetune-CoT paradigm [16], which uses unstructured CoT explanations from the teacher as a supervision signal.
+This distillation approach has been demonstrated to be effective with multi-step mathematical reasoning [5].
+However, relying on unstructured CoT often leads SLMs to learn spurious correlations rather than deep causal features, thereby reducing robustness on complex data [13, 29].
 Consequently, recent research increasingly favors methods that enforce structural decomposition and explicit causality.
-These approaches include creating constrained distillation pipelines, with the aim of removing ambiguous input context through advanced structural formats.
-For example, SocraticCoT provides an explicit structural format by guiding the model through reasoning using defined subquestion-solution pairs [shridhar2023distilling].
-Alternatively, Mixed Distillation integrates the CoT with a formal verifiable Program-of-Thought reasoning to provide a less ambiguous supervision signal [li2023mixed].
+These approaches include creating constrained distillation pipelines with the aim of removing ambiguous input context through advanced structural formats.
+For example, SocraticCoT provides an explicit structural format by guiding the model through reasoning using defined subquestion-solution pairs [29].
+Alternatively, Mixed Distillation integrates the CoT with a formal verifiable Program-of-Thought reasoning to provide a less ambiguous supervision signal [13].
 In the Text-to-SQL domain, structured KD has seen limited uptake.
-Previous studies, while validating the effectiveness of distillation in this area, have generally relied on less structured signals, such as schema-based finetuning to improve schema linking [hong2024next], the inclusion of enterprise-specific custom examples [hoang-etal-2025-distill], and unstructured CoT reasoning transfer [rossiello2025rationalization].
+Previous studies, while validating the effectiveness of distillation in this area, have generally relied on less structured signals, such as schema-based finetuning to improve schema linking [11], the inclusion of enterprise-specific custom examples [10], and unstructured CoT reasoning transfer [27].
 The proposed Struct-SQL framework systematically evaluates whether structure-based reasoning transfer, specifically QP-CoT, improves the distillation for Text-to-SQL.
+5 Limitations
+Struct-SQL significantly reduces syntactic errors through the use of structured reasoning; however, several constraints remain:
+Token Overhead and Efficiency: Struct-SQL significantly reduces syntactic errors through structured reasoning; however, this approach introduces inference overhead.
+The generation of an intermediate query plan requires approximately 3.6 times more tokens than ReasonSQL (Table 3).
+This increase impacts latency and operational costs; nevertheless, these requirements may still be lower than those of teacher LLMs.
+It would be worthwhile to investigate more concise query plan templates to explore this trade-off.
+Performance Bound: The student model’s capabilities are fundamentally capped by the teacher’s fidelity.
+Our distillation pipeline utilizes a strict filtering mechanism, training only on instances where the teacher model generates an execution-correct SQL query.
+Although this quality control is essential to prevent error propagation and provide a stable signal for small models [38], it inherently skews the training data toward problems that the teacher can solve, potentially limiting the student’s exposure to edge cases that the teacher itself cannot solve [25].
+As a result, even with a competitive 60.42% EX on the BIRD test set, the performance remains bound by the teacher’s own limitations in handling highly complex multi-level joins and set operations, as discussed in Figure 4(a).
+To improve performance further, future work could explore using multiple teachers or human-annotated data to cover edge cases that a single source model may miss.
+Benchmark Selection: A fixed reasoning template is used to minimize syntactic hallucinations through a constrained logical path; however, we recognize that it may limit flexibility across highly diverse database dialects.
+We intentionally prioritize evaluation on the BIRD benchmark, as it represents a more challenging dataset than SPIDER.
+As demonstrated in recent literature (e.g., Snowflake Arctic-Text2SQL-R1), state-of-the-art models frequently exceed 88% EX on the SPIDER-test, yet often struggle to surpass 70% on BIRD-dev [34].
+This performance gap justifies our decision to focus exclusively on BIRD.
 6 Conclusion and Future Research
@@
 An additional ablation study demonstrated generalization to a different base model.
-This increased precision, however, comes with a trade-off: the mandatory generation of the intermediate query plan increases the inference token overhead and relies on a fixed reasoning template.
-Additionally, the Student’s performance remains bound by the Teacher’s fidelity, particularly with respect to complex join and set operations.
-Future research should therefore explore integrating or combining this structured reasoning with other established strategies to determine the optimal signal for distillation efficiency.
 Given the demonstrated effectiveness of structured signals in the refinement of syntactic efficiency and logical reasoning, further research is warranted to validate the effect of this structured distillation approach on other complex reasoning tasks beyond Text-to-SQL.
@@
 We thank Khalid Eidoo, Abdul Hamid Dabboussi, María Rodríguez-Liñán and Ting Fung Lam for their insightful technical discussions and valuable feedback throughout the project and during the review of this manuscript.
+Appendix A Structured Reasoning Prompt Template (QP-CoT)
+The QP-CoT prompt, as detailed in Listing 1, comprises a system instruction followed by two few-shot examples that demonstrate the query execution plan methodology.
+The prompt concludes with the target task; for the sake of brevity in this manuscript, the database schema definitions have been abbreviated.
+Appendix B Unstructured Reasoning Prompt Template (CoT)
+The ReasonSQL baseline (Listing 2) utilizes a 4-shot prompt to elicit free-form natural language reasoning.
+By requiring the model to ’think step-by-step’ without a formal execution plan, this baseline represents the standard unstructured distillation signal.
+Appendix C Struct-SQL Output Examples
+This section provides examples of the Struct-SQL model’s outputs across three levels of SQL complexity: Simple, Moderate, and Challenging.
+These examples demonstrate the model’s ability to generate a human-readable Query Plan that mirrors the logical execution steps of a database engine before producing the final SQLite SQL.
+Each entry includes the Question ID and Database ID for traceability.
+The following examples are selected based on the first successful generation encountered for each difficulty tier within the development dataset to ensure an unbiased representation of the model’s capabilities.